LianaT:

Please please help I don't understand this question at all 😭 Devon and his friends bought strawberry wafers for $3 per packet and chocolate wafers for $1 per packet at a carnival and spent a total of $30. They bought total of 22 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of strawberry wafers and the number of packets of chocolate wafers that Keith and his friends bought at the carnival. Define the variables used in the equations. (5 points) Part B: How many packets of chocolate wafers and strawberry wafers did they buy? Explain how you got the answer. (5 points)

LianaT:

Also if this question has already been answered then can someone please tell me where to look, I'm kind of new to this website.

Renne:

It's not a choice.

Renne:

Part A and part B are instructions.

AstridBella:

ok

LianaT:

Yeah I have to type the answer

Renne:

Let me see how I can help.

Renne:

Okay so they spent a total of $30 to buy 22 packets right?

LianaT:

Right but wouldn't that mean there are multiple combinations of 3$ and 1$ to add up to that?

Renne:

Yes it would.

Renne:

But in order to do that you need to know how many times they bought each item which would most likely be 11 and 11 because they bought 22.

Renne:

But it might be different because it's 3 and 1

Renne:

Hold on would you mind if I brought some help here? You know some help to help the helper?

LianaT:

Sure =)

Renne:

Okay I don't want to overwelm him though so first I will ask if he is willing to help

Renne:

Okay. He is willing to help and I will also continue to help in any way I can.

LianaT:

Thank you C=

Renne:

Of course

Renne:

I don't know if my calculations were right in this situation-

Renne:

I have another question. Were you given an example-? Maybe something similar to this? If I can interpret the example then I am pretty sure I can figure this out. Have you done anything similar to this? Because I feel like you may need to add and multiply and maybe divide but I need just a little more info to do so. If not I will work with what I have.

LianaT:

No my teacher gave me a slideshow with examples of the other 3 questions but not this one =\

OLIVER69:

Simple for part A, you will need to make two equations one for the price of the wafers and the other for how many wafers there are. You also need to variable so x and y are the simplest to use. x = strawberry y = chocolate First equation would be 3x + 1y = 30 Second equation would be 3x - 1y = 22 Part B you would solve the equation like so: Rewrite equations: 3x+y=30; 3x−y=22 Step: Solve 3x+y=30 for y: 3x+y=30 3x+y+−3x=30+−3x (Add -3x to both sides) y=−3x+30 Step: Substitute−3x+30 fo ry in 3x−y=22: 3x−y=22 3x− (−3x+30)=22 6x−30=22 (Simplify both sides of the equation) 6x−30+30=22+30 (Add 30 to both sides) 6x=52 6x6=52/6 (Divide both sides by 6) x=26/3 Step: Substitute 26/3 for x in y=−3x+30: y=−3x+30 y=−3(263)+30 y=4 (Simplify both sides of the equation) Answer: x=26/3 and y=4 (makes sense?)

Renne:

...I need to do better-

LianaT:

Oh man thank you both so much it makes a lot more sense now C=